OpenStudy (anonymous):

Nutrition: One eight-ounce glass of apple juice and one eight-ounce glass of orange juice contain a total of 185 milligrams of vitamin C. Two eight-ounce glasses of apple juice and three eight-ounce glasses of orange juice contain a total of 452 milligrams of vitamin C. How much vitamin C is in an eight-ounce glass of each type of juice. Someone please help!

OpenStudy (anonymous):

let x = amount in orange and y = amount in apple x+y = 185 2x + 3y = 452 solve these two and youve got your answer

OpenStudy (anonymous):

Thank you! Do I set the first one up to look like this? 2x=452-3y

OpenStudy (anonymous):

sorry its the other way around x = apple y = orange

OpenStudy (anonymous):

yes but i think easiest way is to multiply first equation by 2 and subtrcat them the elimination method

OpenStudy (anonymous):

2x + 2y = 370 2x + 3y = 452 subtract first from second and u get y = 82 then plug this into x + y = 185 to get x

OpenStudy (anonymous):

x=103

OpenStudy (anonymous):

yes - thtas the apple juice

OpenStudy (anonymous):

always look for easiest methos - substitution method would have been longer here due to the big numbers

OpenStudy (anonymous):

That makes sense. Now for me to get the orange juice I plug x in and solve right?

OpenStudy (anonymous):

- i've done that already when i did the subtract - see above

OpenStudy (anonymous):

i wish my cat wouldnt jump on the keyboerd!!!!! lol

OpenStudy (anonymous):

Oh oops I wasn't paying attention to that I'm sorry! Yeah Lol only mine is a miniature wienie dog.

OpenStudy (anonymous):

dogs are ok but cats are my favourite so u ok with the answer now?

OpenStudy (anonymous):

Yes! Are you good at the back substitution method with fractional numbers? I am good at it until I get to the fractions. It is my last problem.

OpenStudy (anonymous):

i not quite sure what you mean - but try me

OpenStudy (anonymous):

oh really sorry byrd but i gotta go - i forgot to pick up my grandson - hes just phoned

OpenStudy (anonymous):

Okay I am supposed to start with z and insert it using back-substitution.